Jimmy the Rat invested part of $20,000 at 12% annually and the rest at 14%. The total income from these investments is $2560. How much did Jim invest at each rate?
Answer by stanbon(75887)

Jimmy the Rat invested part of $20,000 at 12% annually and the rest at 14%. The total income from these investments is $2560. How much did Jim invest at each rate?
---------
Equation:
interext + interest = 2560
0.12x + 0.14(20,000-x) = 2560
12x + 14*20000 - 14x = 256000
-2x = 256000-14*20000
-2x = -24000
x = $12,000 (amt invested at 12%)
20,000-x = $8,000 (amt invested at 14%)
========================================
Cheers,
Stan H.